OpenStudy (anonymous):

Help please. Evaluate the integral by making the given substitution. The integral of xe^(-x^2/2) dx with u=-x^2/2

OpenStudy (anonymous):

du = -x, so you can rewrite the integral as: \[-\int\limits e^u du\] Do you see how to do it from there?

OpenStudy (anonymous):

I understand that part, but the next step is what i do not understand. Could you work it out please?

OpenStudy (anonymous):

Oh, sorry. After that, you find the antiderivative, which is just e^u, and then add in the negative sign, so -e^u. Then, you know that u=-x^2/2, so it's -e^(-x^2/2)+C.

OpenStudy (anonymous):

Do you see how that works?

OpenStudy (anonymous):

I see. Thank you. Also for the -x^2/2, the 2s can cancel out right? which is why its just -x?

OpenStudy (anonymous):

Right, the 2's cancel when you take the derivative.

OpenStudy (anonymous):

cool cool. thank you so much. you are very helpful.

OpenStudy (anonymous):

No problem.